# Evaluating Time Series Foundation Models on Noisy Periodic Time Series

### v1

@@ 1 Introduction
In recent years, foundation models have ushered in dramatic breakthroughs to the field of machine learning[1], primarily due to their remarkable ability to excel at zero-shot and few-shot learning tasks. While this revolution is the most remarkable in the context of language generation[2], such models have gained prominence in other domains as well, including for time series data. However, while several foundation models for time series (TSFM) forecasting have been presented, rigorous comparative analysis of these models remains scarce, and our understanding of their performance beyond benchmark datasets is still limited. Unlike traditional statistical models, foundation models lack well-defined assumptions and theoretical guarantees and bounds, making it uncertain how they would perform forecasting previously unseen time series with arbitrary statistical properties.
@@ 3.1 Synthetic Data Generation
A wide variety of real-world data can be approximated by an aggregate of sinusoids of varying frequency, mixed with additive noise. These include meteorological data, radio and audio waves, EEG and ECG sequences as well as electromagnetic signals of other kinds. Often, a discrete-time, sampled version of the original continuous signal is available. Because of their relatively stable statistical properties over long time horizons, they make for an interesting case study for multi-step forecasting models. With this motivation, the synthetic data is generated as weighted sums of sinusoids with additive Gaussian noise. First, a continuous-time version is produced as:
@@
where $w(t)$ is zero-mean, uncorrelated Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$. As long as the frequencies $f_{n}$ are rational, the noise-free part of $X(t)$ is periodic. Following the Nyquist theorem, to avoid aliasing, $X(t)$ is uniformly sampled at intervals of $\tau>2f_{\max}$, where $f_{\max}=\max\{f_{n}\}$, to obtain the test time series. We consider two versions of (1) to generate two sets of test data, that we call set A and set B. In set A, the higher frequencies are generated as multiple of the lowest frequency $f_{1}$, so that $f_{n}=nf_{1}$, with $f_{1}$ being an integer. In this case the time series takes the form:
@@
For both sets, we vary the number of components N in [1, 2, 3, 5, 8, 12, 20], and randomly generate 20 sets of amplitudes $A_{n}$ and frequencies $f_{n}$ for each N. The phase $\phi_{n}$ is kept 0. For set A, the principal frequencies $f_{1}$ are randomly chosen, whereas for set B, all the frequencies are randomly picked. For each amplitude-frequency combination, we vary $\sigma^{2}$ so that the signal-to-noise ratio (SNR) in dB takes values of 2, 5, 10, 15, 20 and 30. Finally the sampling rate $\tau$ is also varied as the following multiples of $f_{\max}$: [2.1, 2.5, 3, 5, 10, 20], where $f_{\max}=\max\{f_{n}\}$. With these choices, for sets A and B we generate 5,040 time series each, each with 576 observations. We consider a context length of 512 and a forecast horizon of 64 for each model tested.

### v2

@@ 1 Introduction
In recent years, foundation models have ushered in dramatic breakthroughs to the field of machine learning[1], primarily due to their remarkable ability to excel at zero-shot and few-shot learning tasks. While this revolution is the most remarkable in the context of language generation[2], such models have gained prominence in other domains as well, including for time series data. However, while several foundation models for time series (TSFM) forecasting have been presented, rigorous comparative analysis of these models remains scarce, and our understanding of their performance beyond benchmark datasets is still limited. Unlike traditional statistical foundation models lack well-defined assumptions and theoretical guarantees and bounds, making it uncertain how they would perform forecasting previously unseen time series with arbitrary statistical properties.
@@ 3.1 Synthetic Data Generation
A wide variety of real-world data can be approximated by an aggregate of sinusoids of varying frequency, mixed with additive noise. These include meteorological data, radio and audio waves, EEG and ECG sequences, as well as electromagnetic signals of other kinds. Often a discrete-time, sampled version of the original continuous signal is available. Because of their relatively stable statistical properties over long time horizons, they make for an interesting case study for multi-step forecasting models. With this motivation, the synthetic data is generated as weighted sums of sinusoids with additive Gaussian noise. First, a continuous-time version is produced as:
@@
where $w(t)$ is zero-mean, uncorrelated Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$. As long as the frequencies $f_{n}$ are rational, the noise-free part of $X(t)$ is periodic. Following the Nyquist theorem, to avoid aliasing, $X(t)$ is uniformly sampled at the rate of $f^{*}>2f_{\max}$, where $f_{\max}=\max\{f_{n}\}$, to obtain the test time series. We consider two versions of (1) to generate two sets of test data, that we call set A and set B. In set A, the higher frequencies are generated as multiple of the lowest frequency $f_{1}$, so that $f_{n}=nf_{1}$, with $f_{1}$ being an integer. In this case the time series takes the form:
@@
For both sets, we vary the number of components N in [1, 2, 3, 5, 8, 12, 20], and randomly generate 20 sets of amplitudes $A_{n}$ and frequencies $f_{n}$ for each N. The phase $\phi_{n}$ is kept 0. For set A, the principal frequencies $f_{1}$ are randomly chosen, whereas for set B, all the frequencies are randomly picked. For each amplitude-frequency combination, we vary $\sigma^{2}$ so that the signal-to-noise ratio (SNR) in dB takes values of 2, 5, 10, 15, 20 and 30. Finally the sampling rate $f^{*}$ is also varied as the following multiples of $f_{\max}$: [2.1, 2.5, 3, 5, 10, 20], where $f_{\max}=\max\{f_{n}\}$. With these choices, for sets A and B we generate 5,040 time series each, each with 576 observations. We consider a context length of 512 and a forecast horizon of 64 for each model tested.